# ReAgent: Reversible Multi-Agent Reasoning for Knowledge-Enhanced Multi-Hop QA

### v1

@@ 7 Conclusion
Despite its advantages, our framework inherits several constraints:
  • Complex Coordination: Multi-agent interaction and backtracking mechanisms introduce additional computational overhead and higher implementation complexity.
  • Dependence on Retrieval Quality: As with many QA systems, our accuracy hinges on retrieving relevant and sufficiently detailed evidence. Misleading or scarce information can impair performance, even with backtracking.
  • Scalability to Larger Contexts: While our approach manages conflicts through localized rollback, handling very large or highly diverse corpora may require more sophisticated retrieval and aggregation strategies to maintain efficiency.
  • Model Assumptions: We rely on the underlying language models’ capacity for reliable verification. If the base model’s reasoning is systematically biased, backtracking alone may not fully correct errors.
Ethic Statement
Our method focuses on improving the interpretability and robustness of multi-hop question answering rather than amplifying any content that could be personally or socially harmful. We do not incorporate user-sensitive or private data, and the knowledge sources employed are publicly accessible. However, as with any automated system, inaccurate or biased inputs may lead to flawed inferences if not properly filtered or monitored. We encourage careful data curation and ongoing human oversight to mitigate unintended consequences. Furthermore, we recommend that users verify critical or sensitive outputs, since backtracking techniques reduce—but cannot entirely eliminate—residual errors inherent in large language models.
@@ A.7 Usage and Integration Notes
Appendix B Case Study
B.1 Puzzle Statement
@@
B.2 Agents and Their Roles
@@
B.3 Walkthrough of the Reasoning and Backtracking
B.3.1 Initial Question Decomposition
@@
B.3.2 Retrieval and a Local Conflict
@@
B.3.3 Assembling Partial Answers & Uncovering a Global Conflict
@@
B.3.4 Global Backtracking and Final Resolution
@@
B.4 Key Observations
@@
Appendix C Traditional Multi-hop QA Tasks
1. The Question
C.0.1 User’s question:
To answer this, the system must:
  1. 1.
    Identify the state that hosted the 1984 Summer Olympics.
  2. 2.
  3. 3.
    Confirm the capital city’s population is indeed smaller.
  4. 4.
    Provide the name of that state.
C.0.2 Agents and Their Roles
  1. 1.
    Question-Decomposer Agent ($A_{Q}$)
    Splits the complex question into sub-questions:
    • Sub-question 1: Which U.S. state hosted the 1984 Summer Olympics?
  2. 2.
    Retriever Agent ($A_{R}$)
  3. 3.
    Verifier Agent ($A_{V}$)
  4. 4.
    Answer-Assembler Agent ($A_{A}$)
  5. 5.
    Supervisor Agent ($A_{S}$)
  6. 6.
    Controller Agent ($A_{C}$)
C.0.3 Walkthrough of the Reasoning and Backtracking
3.1 Initial Question Decomposition
  1. 1.
    $A_{Q}$ receives the question:
    $A_{Q}$ parses the main query:
  2. 2.
    $A_{Q}$ produces sub-questions:
    • $q_{1}$: “Which U.S. state hosted the 1984 Summer Olympics?”
    • $q_{4}$: “Return the state if the capital’s population is smaller.”
3.2 Retrieval and a Local Conflict
  1. 1.
    $A_{R}$ retrieves answers:
    • For $q_{1}$, $A_{R}$ returns:
    • For $q_{2}$, $A_{R}$ retrieves partial data:
      • Capital of California is Sacramento.
      • Largest city is Los Angeles.
          Sacramento population: $\sim 508,000$ (from official records).
          Los Angeles population: $\sim 3,900,000$ (correct).
  2. 2.
    $A_{V}$ (Verifier) checks local consistency:
    • Data point $D_{1}$: 508,000
    • Data point $D_{2}$: 1,500,000
    “Population of Sacramento cannot be both 508k and 1.5M.”
  3. 3.
    Local Backtracking by $A_{V}$:
3.3 Assembling Partial Answers & Uncovering a Global Conflict
  1. 1.
    $A_{A}$ (Answer-Assembler) integrates partial conclusions:
    So far, the state that hosted the 1984 Olympics is California.
    The capital is Sacramento (population: $\sim 508,000$).
    The largest city is Los Angeles (population: $\sim 3,900,000$).
  2. 2.
    $A_{A}$ composes a preliminary final:
  3. 3.
    New conflict introduced:
  4. 4.
    $A_{V}$ tries local resolution:
  5. 5.
    Conflict escalates to the Supervisor ($A_{S}$):
    $A_{S}$ identifies the minimal conflicting set:
    • Capital(California, Sacramento)
    • Capital(California, Los Angeles)
3.4 Global Backtracking and Final Resolution
  1. 1.
    Global Backtracking:
  2. 2.
    Controller ($A_{C}$) challenges the suspicious assertion:
    The system confirms that Los Angeles was never the official capital.
  3. 3.
    Assemble the final consistent answer:
    With the conflicting statement removed, the pipeline reaffirms:
    1. (i)
      The 1984 Summer Olympics took place in Los Angeles, California.
    2. (ii)
      California’s capital is Sacramento.
    3. (iii)
    Final Answer: California.
C.0.4 4. Key Observations

### v2

@@ 7 Conclusion
In this paper, we present the ReAgent model for multi-hop QA, which introduces explicit backtracking mechanisms that allow the system to correct errors during reasoning, leading to more accurate and reliable answers. However, this design introduces certain limitations. First, the ability to backtrack multiple times can increase the overall inference time, potentially making the reasoning process less efficient. Second, although ReAgent demonstrates improved performance over selected reasoning models, its more complex architecture may reduce its robustness in scenarios with limited resources or noisy inputs. In the future, we aim to improve the design of the agents, with a particular focus on enhancing their collaboration strategies to further reduce error propagation and improve reasoning efficiency.
@@ A.7 Usage and Integration Notes
Appendix B Analysis of Computational Cost
We conduct a comparative analysis of computational cost on the HotpotQA, 2Wiki, and MuSiQue datasets across GPT-4o, reasoning models O1 and O3, and our ReAgent. The evaluation includes average inference calls, inference time (s), input/output tokens, and total cost(\$), as shown in Table 2. The results highlight that our ReAgent, built on GPT-4o, achieves superior performance compared to reasoning models O1 and O3, while incurring significantly lower computational cost. Although its cost is higher than the GPT-4o baseline, ReAgent delivers substantially better performance. These findings demonstrate the effectiveness of our method in balancing performance and efficiency.
Model Avg. Avg. Avg. Avg. Total
Calls Time(s) Input(T) Output(T) Cost(\$)
GPT-4o 1 2.7 9 289 5  69
O1 1 43 9 289 1 471  645
O3 1 27 9 289 1 471  546
Ours (ReAgent,(GPT-4o)) 29 46 12 400 1 920 275
Table 2: Cost Comparison. T denotes Tokens. Note that the input token counts are aggregated across the three datasets. Since the complete reasoning outputs for O1 and O3 were unavailable, their output token counts are estimated based on the reasoning outputs generated by Qwen-3-235B.
Appendix C Case Study on Puzzle Solving
Algorithm 1 ReAgent Multi-Agent Puzzle Solving
1:Input: Puzzle statements from A, B, C, D.
3:Initialization
[…]
6:Round 1: Checking A as culprit
10:Round 2: Checking B as culprit
13:$T_{6}:$ Conflict Detector sees consistency, no rollback needed.
14:Rounds 3 and 4: Checking C or D as culprit
We illustrate our multi-agent backtracking (ReAgent) through a single-culprit puzzle with four suspects {A, B, C, D}. Each suspect gives statements about who might be guilty or lying. The puzzle’s only correct solution is that B is the culprit, but identifying this requires partially retracting initial assumptions along the way, as shown in Algorithm 1.
Why a Rollback Mechanism is Needed. Naive single-pass (or single-thread) models often fixate prematurely on one suspect, disregard contradictory evidence, and produce unsound conclusions. By contrast, our multi-agent approach enumerates possible culprits in parallel, detects conflicts, and rolls back to revise incorrect assumptions.
@@ Appendix C Case Study on Puzzle Solving
Appendix D Case Study on Traditional Multi-hop QA Tasks
D.1 User’s question:
To answer this, the system must:
  1. 1.
    Identify the state that hosted the 1984 Summer Olympics.
  2. 2.
  3. 3.
    Confirm the capital city’s population is indeed smaller.
  4. 4.
    Provide the name of that state.
D.2 Agents and Their Roles
@@
D.3 Walkthrough of the Reasoning and Backtracking
Initial Question Decomposition
@@ D.3 Walkthrough of the Reasoning and Backtracking
Retrieval and a Local Conflict
@@
Assembling Partial Answers & Uncovering a Global Conflict
@@
Global Backtracking and Final Resolution
@@
D.4 Full Comparison with Baseline Models (Figure 7)
Figure 7: Full Comparison with Baseline Models
D.5 Key Observations
@@
Algorithm 2 Multi-Agent Reversible Reasoning
1:$Q_{0}$: Main question
$A_{Q}.\mathrm{decompose}(Q_{0})\to\{q_{1},q_{2},\dots\}$
3:Broadcast $\mathrm{assert}(q_{i})$ to $A_{R},A_{V},A_{A}$
4:for each $q_{i}$ in $\{q_{1},q_{2},\dots\}$ concurrently do
5:     $E\leftarrow A_{R}.\mathrm{retrieve}(q_{i})$
6:     $A_{V}.\mathrm{verify}(E)$
7:     if $A_{V}$ detects conflict locally then
8:         $A_{V}.\mathrm{BacktrackLocal}(r_{V})$
9:         if conflict persists then
10:              raise Conflict to $A_{S}$
11:         end if
12:     end if
13:     $A_{A}.\mathrm{storePartialAnswer}(q_{i},E)$
14:end for
15:if Global conflict signaled then
16:     $A_{S}.\mathrm{HolisticUpdate}(\Pi_{j})$
17:     if $A_{C}$ intervention needed then
18:         $A_{C}.\mathrm{challenge}(\varphi)$ or override
19:     end if
20:end if
22:return Final Answer
@@
Appendix E Reversible Reasoning Algorithm